# Strategic Informed Trading and the Value of Private Information

### v1

@@ Appendix B Proofs from Section 1 and Appendix A
$\begin{split}p(H)&=p_{0}+\frac{\beta+\lambda_{I}}{\lambda_{U}R+\beta+\lambda_{I}}\left((1-R)X+\sqrt{R(1-R)}P_{X}^{-1/2}\mathcal{E}_{I}+\frac{1}{\alpha_{I}}RP_{X}^{-1}Z_{N}-(1-R)p_{0}\right).\end{split}$
In (68) note that $p_{I}\to 0$ corresponds to $R\to 1,\beta\to 0$. Therefore, almost surely
$\begin{split}\lim_{p_{I}\to 0}p(H)&=p_{0}+\frac{\lambda_{I}}{(\lambda_{U}+\lambda_{I})}P_{X}^{-1}\frac{Z_{N}}{\alpha_{I}}=p_{0}+\lambda_{I}P_{X}^{-1}\frac{Z_{N}}{\alpha_{I}},\end{split}$
@@
because $\lambda_{U}+\lambda_{I}=1$. This gives the pricing formula in (63) as $\lambda=\lambda_{I}$. As for the positions, first note from (72) below that
@@
$\begin{split}P_{U}&=\frac{(1-R)\beta}{R(1+\beta)}P_{X}\to 0;\qquad P_{I}G=\frac{1-R}{R}P_{X}\left(X+\sqrt{\frac{R}{1-R}}P_{X}^{-1/2}{\mathcal{E}}_{I}\right)\to 0,\\ P_{U}H&=\frac{(1-R)\beta}{R(1+\beta)}P_{X}\left(X+\sqrt{\frac{R}{1-R}}P_{X}^{-1/2}\mathcal{E}_{I}+\frac{R}{\alpha_{I}(1-R)}P_{X}^{-1}Z_{N}\right)\to 0.\end{split}$
(63) follows from (55) and (62). We next consider the price impact case. From (68) and (69) we see that $p_{I}\to 0$ additionally implies $(1-R)/p_{I}\to 1$ and $\widehat{y}\to\lambda_{I}/\lambda_{U}$. From (45), (51), and (73)
$\begin{split}p_{\iota}(H_{\iota})&=p_{0}+\frac{\widehat{y}}{1+2\widehat{y}}\times\left((1-R)X+\sqrt{R(1-R)}P_{X}^{-1/2}\mathcal{E}_{I}+\frac{R(1+\widehat{y})}{\alpha_{I}}P_{X}^{-1}Z_{N}-(1-R)p_{0}\right)\\ &\qquad+\frac{\lambda_{I}\widehat{y}\left(\beta+(1+\widehat{y})^{2}\right)}{\lambda_{U}(1+2\widehat{y})(1+\widehat{y})^{2}}\times P_{X}^{-1}\left(\psi_{I,0}-\widehat{\Pi}\right).\end{split}$
Therefore, almost surely, and using $\lambda_{U}+\lambda_{I}=1$
@@
$\begin{split}\lim_{p_{I}\to 0}p_{\iota}(H_{\iota})&=p_{0}+\frac{\lambda_{I}(\lambda_{U}+\lambda_{I})}{(\lambda_{U}+2\lambda_{I})\lambda_{U}}P_{X}^{-1}\frac{Z_{N}}{\alpha_{I}}+\frac{\lambda_{I}^{2}}{\lambda_{U}(\lambda_{U}+2\lambda_{I})}P_{X}^{-1}\left(\psi_{I,0}-\widehat{\Pi}\right),\\ &=p_{0}+\frac{\lambda_{I}}{1-\lambda^{2}_{I}}P_{X}^{-1}\frac{Z_{N}}{\alpha_{I}}+\frac{\lambda_{I}^{2}}{1-\lambda^{2}_{I}}P_{X}^{-1}\left(\psi_{I,0}-\widehat{\Pi}\right).\end{split}$
As for $I$’s optimal policy, from (6)
$\widehat{\psi}(G,Z_{N})=\psi_{I,0}-\frac{Z_{N}}{\alpha_{I}}+M^{-1}(p_{\iota}(H_{\iota})-V).$
From (59) and (68) we find that $M^{-1}\to(\lambda_{U}/\lambda_{I})P_{X}$, and we have just computed the limit of $p_{\iota}(H_{\iota})$. As for $V$, from (57) as well as (74) and (75) below we deduce
(71) $\begin{split}V&=p_{0}+\frac{\beta+\lambda_{I}(1+\widehat{y})^{2}}{\lambda_{U}(1+\widehat{y})^{2}}P_{X}^{-1}(\psi_{I,0}-\widehat{\Pi}).\end{split}$
Therefore, $V\to p_{0}+(\lambda_{I}/\lambda_{U})P_{X}^{-1}(\psi_{I,0}-\widehat{\Pi})$. Putting everything together
$\begin{split}\lim_{p_{I}\to 0}\widehat{\psi}_{I,\iota}(G,Z_{N})&=\psi_{I,0}-\frac{Z_{N}}{\alpha_{I}}+\frac{\lambda_{U}}{\lambda_{I}}P_{X}\bigg{(}p_{0}+\frac{\lambda_{I}}{1-\lambda^{2}_{I}}P_{X}^{-1}\frac{Z_{N}}{\alpha_{I}}+\frac{\lambda_{I}^{2}}{1-\lambda_{I}^{2}}P_{X}^{-1}\left(\psi_{I,0}-\widehat{\Pi}\right)\\ &\qquad\qquad-p_{0}-\frac{\lambda_{I}}{\lambda_{U}}P_{X}^{-1}(\psi_{I,0}-\widehat{\Pi})\bigg{)},\\ &=\psi_{I,0}-\left(1-\frac{\lambda_{U}}{1-\lambda^{2}_{I}}\right)\frac{Z_{N}}{\alpha_{I}}-\left(1-\frac{\lambda_{I}\lambda_{U}}{1-\lambda_{I}^{2}}\right)(\psi_{I,0}-\widehat{\Pi}),\\ &=\psi_{I,0}-\frac{\lambda_{I}}{1+\lambda_{I}}\frac{Z_{N}}{\alpha_{I}}-\frac{1}{1+\lambda_{I}}\left(\psi_{I,0}-\widehat{\Pi}\right).\end{split}$
The result for $\widehat{\psi}_{ns,U,\iota}$ can be deduced from the clearing condition (4), which using the current notation implies that in the limit
@@
$\widehat{\psi}_{ns,U,\iota}=\frac{1}{1-\lambda_{I}}\left(\widehat{\Pi}-\lambda_{I}\frac{Z_{N}}{\alpha_{I}}-\lambda_{I}\widehat{\psi}_{ns,I,\iota}\right).$
Appendix C Proofs from Sections 2 & 4
@@
The first statement follows directly from (30) and $\widehat{y}>0$. For the second, direct calculation shows for $p_{U},p_{U,\iota}$ from (29) and using the notation in (68) that
(72) $\begin{split}p_{U}&=\frac{1-R}{R}\frac{\beta}{1+\beta};\qquad p_{U,\iota}=\frac{1-R}{R}\left(\frac{\beta}{(1+\widehat{y})^{2}+\beta}\right).\end{split}$
Thus, $p_{U}>p_{U,\iota}$ is equivalent to $1<(1+\widehat{y})^{2}$, which holds as $\widehat{y}>0$. ∎
We next prove Proposition 2.3 in the general setting of Appendix A. Proposition 2.3 will follow as in Section 1 we have (see (19)) $p_{0}=\psi_{I,0}=\widehat{\Pi}$.
Proposition C.1.
Let Assumption A.4 hold and recall (55). The pricing functions from (58) and (61) specify to
(73) $\begin{split}p_{\iota}(h_{\iota})&=p_{0}+\frac{p_{I}\widehat{y}}{(1+p_{I})(1+2\widehat{y})}\times\left(h_{\iota}-p_{0}\right)+\frac{\lambda\widehat{y}\left(\kappa p_{I}+(1+\widehat{y})^{2}\right)}{(1-\lambda)(1+2\widehat{y})(1+\widehat{y})^{2}}\times P_{X}^{-1}\left(\psi_{I,0}-\widehat{\Pi}\right),\\ p(h)&=p_{0}+\frac{p_{I}(\kappa p_{I}+\lambda)}{1-\lambda+(1+p_{I})(\kappa p_{I}+\lambda)}\times\left(h-p_{0}\right).\end{split}$
Proof of Proposition C.1.
We start with $p(h)$ from (61). From (61), (72) and $\lambda_{I}+\lambda_{U}=1$ we obtain
$\begin{split}P=\frac{(1-R)(\beta+\lambda_{I})}{R(1+\beta)}P_{X};\quad P+P_{X}=\frac{\lambda_{U}R+\beta+\lambda_{I}}{R(1+\beta)}P_{X};\quad(P_{X}+P)^{-1}P=\frac{(1-R)(\beta+\lambda_{I})}{\lambda_{U}R+\beta+\lambda_{I}}1_{d},\end{split}$
and hence the formula for $p(h)$. We next consider $p_{\iota}(h)$ from (58) and drop the explicit dependence on $\widehat{y}$ in $M,{\mathcal{M}},\mathcal{V},V$. First, (59) implies $M{\mathcal{M}}=(1-R)\widehat{y}/(1+2\widehat{y})1_{d}$ which gives the $h_{\iota}-p_{0}$ term. As for the $\psi_{I,0}-\widehat{\Pi}$ term, using (57), (72) we find
(74) $P_{\iota}=\frac{(1-R)(\beta+\lambda_{I}(1+\widehat{y})^{2})}{R(\beta+(1+\widehat{y})^{2})}P_{X}.$
This implies
$\begin{split}P_{\iota}^{-1}+P_{X}^{-1}&=\frac{\beta+(\lambda_{U}R+\lambda_{I})(1+\widehat{y})^{2}}{(1-R)(\beta+\lambda_{I}(1+\widehat{y})^{2})}P_{X}^{-1},\\ P_{X}^{-1}-M{\mathcal{M}}\left(P_{\iota}^{-1}+P_{X}^{-1}\right)&=\frac{(1+\widehat{y})\left(\beta+\lambda_{I}(1+\widehat{y})^{2}-\lambda_{U}R\widehat{y}(1+\widehat{y})\right)}{(1+2\widehat{y})(\beta+\lambda_{I}(1+\widehat{y})^{2})}P_{X}^{-1},\end{split}$
as well as
(75) $P_{I}-P_{\iota}=\frac{(1-R)\lambda_{U}(1+\widehat{y})^{2}}{R(\beta+(1+\widehat{y})^{2})}P_{X}.$
Therefore,
$\begin{split}&\left(P_{X}^{-1}-M{\mathcal{M}}\left(P_{\iota}^{-1}+P_{X}^{-1}\right)\right)P_{\iota}(P_{I}-P_{\iota})^{-1}=\frac{\beta+\lambda_{I}(1+\widehat{y})^{2}-\lambda_{U}R\widehat{y}(1+\widehat{y})}{\lambda_{U}(1+2\widehat{y})(1+\widehat{y})}.\\ \end{split}$
(69) implies $\lambda_{U}R\widehat{y}(1+\widehat{y})=\lambda_{I}(1+\widehat{y})+\lambda_{U}\beta+\lambda_{I}\beta/(1+\widehat{y})$ and hence
$\beta+\lambda_{I}(1+\widehat{y})^{2}-\lambda_{U}R\widehat{y}(1+\widehat{y})=\frac{\lambda_{I}\widehat{y}(\beta+(1+\widehat{y})^{2})}{1+\widehat{y}},$
which gives the result.
@@
We retain the notation in (68) and suppress $\widehat{y}$ from the $M,{\mathcal{M}},\mathcal{V},V$ functions. First, direct calculation using (32) shows
$\frac{2}{1-R}\left(m_{g}-m_{g,\iota}\right)=\frac{1}{1+2\widehat{y}}+\frac{\beta+\lambda_{I}-\lambda_{U}R}{\beta+\lambda_{I}+\lambda_{U}R}.$
This gives the result when $\beta+\lambda_{I}>\lambda_{U}R$. When $\lambda_{U}R>\beta+\lambda_{I}$ define $\widetilde{y}$ through
$\frac{1}{1+2\widetilde{y}}=\frac{\lambda_{U}R-(\beta+\lambda_{I})}{\lambda_{U}R+\beta+\lambda_{I}}.$
Thus, $m_{g}>m_{g,\iota}$ if and only if $\widehat{y}<\widetilde{y}$. In the proof of Proposition 1.5 we showed if we define $g$ by the right side of (18) (see also (69)) , then $g$ is strictly decreasing at $\widehat{y}$. Thus, if $g(\widetilde{y})<0$ it must be that $\widehat{y}<\widetilde{y}$. Indeed, $\widehat{y}=\widetilde{y}$ is not possible, and if $\widehat{y}>\widetilde{y}$ then there must be some $\check{y}>0$ with $g(\check{y})=0$, but by the uniqueness statement in Proposition 1.5 we know this is not possible as well.
@@
$\begin{split}g(\widetilde{y})&=\frac{p^{2}}{(p-q)^{2}}\left(1-\frac{pq}{\lambda_{I}(p-q)}\right)+\frac{\lambda_{U}p\beta}{\lambda_{I}(p-q)}+\beta.\end{split}$
As the common denominator $\lambda_{I}(p-q)^{3}$ is positive, we need only show the numerator is negative. The numerator is
$\lambda_{I}p^{2}(p-q)-p^{3}q+\lambda_{U}p\beta(p-q)^{2}+\beta\lambda_{I}(p-q)^{3}.$
If we group terms by powers of $p$ the cubic terms vanish, leaving
@@
We again retain the notation in (68) and suppress $\widehat{y}$ from the $M,{\mathcal{M}},\mathcal{V},V$ functions. Using (61), (59) and (72) calculation shows
$m_{\widehat{\chi}}=\frac{\beta+\lambda_{I}}{\lambda_{U}};\qquad m_{\widehat{\chi},\iota}=R\widehat{y}.$
Therefore, $m_{\widehat{\chi}}>m_{\widehat{\chi},\iota}$ if and only if $\widehat{y}<\widetilde{y}:=(\beta+\lambda_{I})/(\lambda_{U}R)$. We will show $g(\widetilde{y})<0$ for $g$ defined by the right side of (69), and this will give the result, as the proof of Proposition 2.4 showed for $y>0$ that $g(y)<0$ if and only if $y>\widehat{y}$. To this end, from (69)
$\begin{split}g(\widetilde{y})&=\frac{(\lambda_{U}R+\beta+\lambda_{I})^{2}}{\lambda_{U}^{2}R^{2}}\left(1-\frac{\beta+\lambda_{I}}{\lambda_{I}}\right)+\frac{\beta(\lambda_{U}R+\beta+\lambda_{I})}{R\lambda_{I}}+\beta,\\ &=\frac{-\beta}{\lambda_{I}\lambda_{U}^{2}R^{2}}\left((\lambda_{U}R+\beta+\lambda_{I})^{2}-\lambda_{U}^{2}R(\lambda_{U}R+\beta+\lambda_{I})-\lambda_{I}\lambda_{U}^{2}R^{2}\right),\\ &=\frac{-\beta}{\lambda_{I}\lambda_{U}^{2}R^{2}}\left(\lambda_{U}R(\beta+\lambda_{I})(1+\lambda_{I})+(\beta+\lambda_{I})^{2}\right),\end{split}$
where we have used that $\lambda_{I}+\lambda_{U}=1$. The result follows as $\beta,\lambda_{I},\lambda_{U},R>0$. ∎
@@
The expected value in (44) is negative if and only if
$\widehat{y}>\overline{y}:=\frac{(1+p_{I})(\lambda+\kappa p_{I})}{2(1-\lambda)}.$
Next, as shown in the proof of Proposition 2.4, if we define the function $g$ by the right side of (18) then $\widehat{y}>\overline{y}$ if and only if $g(\overline{y})>0$. Calculations show
$g(\overline{y})=\frac{(\lambda-\kappa p_{I})\left((\lambda+\kappa p_{I})(1+p_{I})+2(1-\lambda)\right)^{2}}{8(1-\lambda)^{2}\lambda}+\frac{\kappa p_{I}((1+p_{I})(\lambda+\kappa p_{I})+2)}{2\lambda}.$
The sign of the right hand side above coincides with that of
$\begin{split}&(\lambda-\kappa p_{I})\left((\lambda+\kappa p_{I})(1+p_{I})+2(1-\lambda)\right)^{2}+4(1-\lambda)^{2}\kappa p_{I}((1+p_{I})(\lambda+\kappa p_{I})+2)\\ &\qquad=(\lambda+\kappa p_{I})\left(((1+p_{I})\lambda+2(1-\lambda))^{2}-4\lambda(1-\lambda)\kappa p_{I}(1+p_{I})-\kappa^{2}p_{I}^{2}(1+p_{I})^{2}\right).\end{split}$
Therefore $\widehat{y}>\overline{y}$ if and only if
$\kappa<\overline{\kappa}:=\frac{1}{p_{I}(1+p_{I})}\left(\sqrt{\left(2\lambda(1-\lambda)\right)^{2}+\left(\lambda(1+p_{I})+2(1-\lambda)\right)^{2}}-2\lambda(1-\lambda)\right).$
Using (17) we see that as the insider becomes very risk averse ($\alpha_{I}\to 0$) we have $\lambda\to 0$ and hence $\overline{\kappa}\to 2/(p_{I}(1+p_{I}))$ while $\kappa\to 0$. Thus $\widehat{y}>\overline{y}$. On the other hand, as the insider becomes sufficiently risk tolerant ($\alpha_{I}\to\infty$), $\overline{\kappa}\to 1/p_{I}$ but $\kappa\to\infty$. Thus, in this instance $\widehat{y}<\overline{y}$.
Appendix D Welfare formulas used in Section 3
We prove results for the general setup in Appendix A and use the notation (68) throughout. We begin by identifying interim certainty equivalents in the price taking case. From (21), the insider has certainty equivalent function (recall (60) implies $\sigma(G,H)=\sigma(G,Z_{N})$)

### v2

@@ B.2. Proofs from Appendix A
$p(H)=p_{{\mathbb{Q}}_{0}}+\frac{(1-R)(\lambda_{I}+\beta)}{\lambda_{I}+\beta+R\lambda_{U}}\left(X_{0}+\sqrt{\frac{R}{1-R}}P_{X}^{-1/2}\mathcal{E}_{I}+\frac{R}{\alpha_{I}(1-R)}P_{X}^{-1}Z_{N}\right).$
From (77) we see that $p_{I}\to 0$ implies $R\to 1,\beta\to 0$. Therefore, almost surely
$\begin{split}\lim_{p_{I}\to 0}p(H)&=p_{{\mathbb{Q}}_{0}}+\frac{\lambda_{I}}{(\lambda_{I}+\lambda_{U})}P_{X}^{-1}\frac{Z_{N}}{\alpha_{I}}=p_{{\mathbb{Q}}_{0}}+\lambda_{I}P_{X}^{-1}\frac{Z_{N}}{\alpha_{I}},\end{split}$
@@
because $\lambda_{U}+\lambda_{I}=1$. This gives the pricing formula in (70) as $\lambda=\lambda_{I}$. As for the positions, from (68), (79) and (81) we find
@@
$\begin{split}\widehat{\psi}_{I}&=\widehat{\Pi}+\frac{1-R}{R}P_{X}\left(X_{0}+\sqrt{\frac{R}{1-R}}P_{X}^{-1/2}\mathcal{E}_{I}\right)-\frac{1}{R}P_{X}\left(p(H)-p_{{\mathbb{Q}}_{0}}\right),\\ \widehat{\psi}_{U}&=\widehat{\Pi}+\frac{(1-R)\beta}{R(1+\beta)}P_{X}\left(X_{0}+\sqrt{\frac{R}{1-R}}P_{X}^{-1/2}\mathcal{E}_{I}+\frac{R}{\alpha_{I}(1-R)}P_{X}^{-1}Z_{N}\right)\\ &\qquad-\left(1+\frac{(1-R)\beta}{R(1+\beta)}\right)\left(p(H)-p_{{\mathbb{Q}}_{0}}\right).\end{split}$
Thus, almost surely as $p_{I}\to 0$ we have
@@
$\begin{split}\widehat{\psi}_{I}&\to\widehat{\Pi}-P_{X}\left(\lambda_{I}P_{X}^{-1}\frac{Z_{N}}{\alpha_{I}}\right)=\widehat{\Pi}-\lambda_{I}\frac{Z_{N}}{\alpha_{I}},\\ \widehat{\psi}_{U}&\to\widehat{\Pi}-P_{X}\left(\lambda_{I}P_{X}^{-1}\frac{Z_{N}}{\alpha_{I}}\right)=\widehat{\Pi}-\lambda_{I}\frac{Z_{N}}{\alpha_{I}}.\end{split}$
We next consider the price impact equilibrium. First, from (52), (62) and (81) we find
(83) $H_{0,\iota}=X_{0}+\sqrt{\frac{R}{1-R}}P_{X}^{-1/2}\mathcal{E}_{I}+\frac{(1+\widehat{y})R}{\alpha_{I}(1-R)}P_{X}^{-1}Z_{N}.$
From (60) and (62) we obtain
@@
$p_{\iota}(H_{\iota})=p_{{\mathbb{Q}}_{0}}+\frac{\widehat{y}(1-R)}{1+2\widehat{y}}\left(X_{0}+\sqrt{\frac{R}{1-R}}P_{X}^{-1/2}\mathcal{E}_{I}+\frac{(1+\widehat{y})R}{\alpha_{I}(1-R)}P_{X}^{-1}Z_{N}.\right).$
From (77) and (78) we see that $p_{I}\to 0$ additionally implies $\widehat{y}\to\lambda_{I}/\lambda_{U}$. Therefore, almost surely, and using $\lambda_{U}+\lambda_{I}=1$ we find
$\lim_{p_{I}\to 0}p_{\iota}(H_{\iota})=p_{{\mathbb{Q}}_{0}}+\frac{\lambda_{I}}{1-\lambda^{2}_{I}}P_{X}^{-1}\frac{Z_{N}}{\alpha_{I}}.$
As for the trading strategies, from (47), (61), (62) and (79) e find
$\begin{split}\widehat{\psi}_{I,\iota}&=\widehat{\Pi}+\frac{1}{1+\widehat{y}}P_{X}p_{{\mathbb{Q}}_{0}}+\frac{1-R}{R(1+\widehat{y})}P_{X}\left(X_{0}+\sqrt{\frac{R}{1-R}}P_{X}^{-1/2}\mathcal{E}_{I}+p_{{\mathbb{Q}}_{0}}\right)-\frac{R}{1+\widehat{y}}P_{X}p_{\iota}(H_{\iota}),\\ \widehat{\psi}_{U,\iota}&=\widehat{\Pi}+P_{X}p_{{\mathbb{Q}}_{0}}+\frac{(1-R)\beta}{R(\beta+(1+\widehat{y})^{2}}P_{X}\bigg{(}X_{0}+\sqrt{\frac{R}{1-R}}P_{X}^{-1/2}\mathcal{E}_{I}\\ &\qquad+\frac{(1+\widehat{y})R}{\alpha_{I}(1-R)}P_{X}^{-1}Z_{N}+p_{{\mathbb{Q}}_{0}}\bigg{)}-\frac{\beta+R(1+\widehat{y})^{2}}{R(\beta+(1+\widehat{y})^{2}}P_{X}p_{\iota}(H_{\iota}).\end{split}$
This gives the almost sure limits as $p_{I}\to 0$
$\begin{split}\widehat{\psi}_{I,\iota}&\to\widehat{\Pi}-\frac{\lambda_{I}}{1+\lambda_{I}}\frac{Z_{N}}{\alpha_{I}};\qquad\widehat{\psi}_{U,\iota}\to\widehat{\Pi}-\frac{\lambda_{I}}{1-\lambda^{2}_{I}}\frac{Z_{N}}{\alpha_{I}}\end{split}$
B.3. Proofs from Sections 2 & 4
As in Section B.2, throughout all proofs of this section we omit the $p,\iota$ subscript from $M_{p,\iota},V_{p,\iota}$ in the price response function $p_{\iota}$ from (5).
@@ B.3. Proofs from Sections 2 & 4
The first statement follows directly from (23) and $\widehat{y}>0$. For the second, using (79) with $P_{X}=1$ we see that $p_{U}>p_{U,\iota}$ is equivalent to $1<(1+\widehat{y})^{2}$, which holds as $\widehat{y}>0$. ∎
@@
The formula for $p_{\iota}$ in (24) is taken directly from (11). As for $p$, using (17) and (80) we find
(84) $p(h)=p_{{\mathbb{Q}}_{0}}+\frac{(1-R)(\lambda_{I}+\beta)}{\lambda_{I}+\beta+R\lambda_{U}}(h-p_{{\mathbb{Q}}_{0}}).$
The expression in (11) follows by the identifications in (77) and (7). ∎
Proof of Proposition 2.3.
We retain the notation in (77) and suppress $\widehat{y}$ from the $M,{\mathcal{M}}$ functions. First, using (25) we readily get that $2\left(m_{g}-m_{g,\iota}\right)/(1-R)=1/(1+2\widehat{y})+(\beta+\lambda_{I}-\lambda_{U}R)/(\beta+\lambda_{I}+\lambda_{U}R)$. This gives the result when $\beta+\lambda_{I}>\lambda_{U}R$. When $\lambda_{U}R>\beta+\lambda_{I}$, define $\widetilde{y}$ through $1/(1+2\widetilde{y})=(\lambda_{U}R-(\beta+\lambda_{I}))/(\lambda_{U}R+\beta+\lambda_{I})$. Thus, $m_{g}>m_{g,\iota}$ if and only if $\widehat{y}<\widetilde{y}$. In the proof of Proposition 1.1 we showed if we define $g$ by the right side of (6) (see also (78)), then $g$ is strictly decreasing at $\widehat{y}$. Thus, if $g(\widetilde{y})<0$ it must be that $\widehat{y}<\widetilde{y}$. Indeed, $\widehat{y}=\widetilde{y}$ is not possible, and if $\widehat{y}>\widetilde{y}$ then there must be some $\check{y}>0$ with $g(\check{y})=0$, but by the uniqueness statement in Proposition 1.1 we know this is not possible as well.
@@
$g(\widetilde{y})=\frac{p^{2}}{(p-q)^{2}}\left(1-\frac{pq}{\lambda_{I}(p-q)}\right)+\frac{\lambda_{U}p\beta}{\lambda_{I}(p-q)}+\beta.$
As the common denominator $\lambda_{I}(p-q)^{3}$ is positive, we need only show the numerator is negative. The numerator is $\lambda_{I}p^{2}(p-q)-p^{3}q+\lambda_{U}p\beta(p-q)^{2}+\beta\lambda_{I}(p-q)^{3}$. If we group terms by powers of $p$ the cubic terms vanish, leaving
@@
We again suppress $\widehat{y}$. Using (77) we have $m_{\widehat{\chi}}=(\beta+\lambda_{I})/\lambda_{U}$ and $m_{\widehat{\chi},\iota}=R\widehat{y}.$ Therefore, $m_{\widehat{\chi}}>m_{\widehat{\chi},\iota}$ if and only if $\widehat{y}<\widetilde{y}:=(\beta+\lambda_{I})/(\lambda_{U}R)$. We will show $g(\widetilde{y})<0$ for $g$ defined by the right side of (78), and this will give the result, as the proof of Proposition 2.3 showed for $y>0$ that $g(y)<0$ if and only if $y>\widehat{y}$. To this end, from (78), $g(\widetilde{y})=(\lambda_{U}R+\beta+\lambda_{I})^{2}\left(1-\beta+\lambda_{I}/\lambda_{I}\right)/(\lambda_{U}^{2}R^{2})+\beta(\lambda_{U}R+\beta+\lambda_{I})/(R\lambda_{I})+\beta$, which equal to $-\beta\left[\lambda_{U}R(\beta+\lambda_{I})(1+\lambda_{I})+(\beta+\lambda_{I})^{2}\right]/(\lambda_{I}\lambda_{U}^{2}R^{2})$, where we have used that $\lambda_{I}+\lambda_{U}=1$. The result follows as $\beta,\lambda_{I},\lambda_{U},R>0$. ∎
@@
Using (77), the expected value in (37) is negative if and only if $\widehat{y}>\overline{y}:=(\lambda_{I}+\beta)/(2R\lambda_{U})$. Next, as shown in the proof of Propositions 2.3, 2.4, if we define function $g$ by the right side of (78) then $\widehat{y}>\overline{y}$ if and only if $g(\overline{y})>0$. Direct calculation yields $g(\overline{y})=\beta[(\lambda_{I}/\beta-1)(2R\lambda_{U}+\lambda_{I}+\beta)^{2}/(8\lambda_{I}\lambda_{U}^{2}R^{2})$ +$4\beta R\lambda_{U}^{2}(\lambda_{I}+\beta+2R)]$. Now, let $\alpha_{I}\to 0$ while keeping everything else fixed. From (77) we see $\lambda_{I}/\beta\to\infty$, $\lambda_{I}\to 0$, $\lambda_{U}\to 1$ and $R$ is fixed. This implies that $g(\overline{y})>0$ and hence $\widehat{y}>\overline{y}$. Conversely, when $\alpha_{I}\to\infty$ we see $\lambda_{I}/\beta\to 0$, $\beta\to\infty$, $\lambda_{I}\to 1$, $\lambda_{U}\to 0$ and $R$ is fixed. This implies that $g(\overline{y})<0$ and hence $\widehat{y}<\overline{y}$. ∎
B.4. Utility formulas used in Section 3
We prove utility results for the setup in Appendix A, under Assumption A.4 and using the notation in (77). We will start with the price impact case, and then proceed to the price taking case. Throughout we recall the initial endowments in (1) and recall that $\widehat{{\mathcal{Y}}}=\widehat{y}1_{d}$ for $\widehat{y}$ from Proposition 1.1. We also use the conditional laws (similar to (71), (74)) for $k\in\left\{\iota,\ \right\}$